DIFFERENTIAL GEOMETRY/PDE SEMINAR

Thursday, October 30, 2003 Padelford C-36 1:30 pm

Quasi-local mass and static metric extension in general relativity

Pengzi Miao (MSRI)

In this talk we study problems motivated by Bartnik’s quasi-local mass definition in general relativity. We first establish a positive mass theorem for a class of piecewise smooth asymptotically flat metrics. Then we investigate the variational effect of boundary mean curvature on the mass and derive Bartnik’s geometric boundary condition on minimal mass extensions of a quasi-convex domain. In the end we discuss the existence of a scalar flat static metric satisfying the derived minimal mass boundary condition.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.